Evaluate the following
f(x) = 6x + 5
f(12) =
f(–1/2) =
f(0) =


Answer by JBarnum(2146)

all this means is simply the x changes values as given
f(x) = 6x + 5
f(12) = 6(12)+5 = 77
f(–1/2) = 6(-1/2)+5 = 2
f(0) = 6(0)+5 = 5